OpenStudy (anonymous):

Carousel Toys has Romper Buckaroos,wooden rocking horses for toddlers,on a 30% markdown sale for $72.09. what was the original price before they were marked down?

OpenStudy (anonymous):

what was the percentage before the markdown sale

OpenStudy (anonymous):

30

OpenStudy (anonymous):

Let the original price be x dollars. hence : \[\huge \frac{30}{100} \times x = 72.09 \] \[\huge 30 \times x = 72.09 \times 100\] \[\huge x = \frac{7209 }{30} = 240.3\] Hence the original price was 240.3 dollars @shanise_92